OpenStudy (anonymous):

Solve for x: log(x+2) + log(x-1)=1 I have no idea how to get to the answer o.O I know the answer is 3, but I can't reach that. I know to use the logarithmic properties to make it log[(x+2)(x-1)]=1 and then just (x+2)(x-1)=10 after that, I got either the answers x=11,8 or x=12,11 What am I doing wrong?

OpenStudy (anonymous):

@SarahEZZMcK

OpenStudy (amistre64):

you can start by baseing each side .. assuming log is a base 10 in your class

OpenStudy (anonymous):

Right, I would've explained if the base was any different :)

OpenStudy (amistre64):

in higher classes, there is only one log, the natural. log(x+2) + log(x-1)=1 10^( log(x+2) + log(x-1)=1) 10^log(x+2) 10^log(x-1) = 10^1 (x+2)(x-1) = 10

OpenStudy (amistre64):

its a quadratic now; we can expand and run the quadratic formula

OpenStudy (amistre64):

x^2 +x - 2 = 10 x^2 +x - 12 = 0

OpenStudy (anonymous):

Why could I not simply take it as x+2=10 and x-1=10?

OpenStudy (anonymous):

Is there a tell for when to break it up like that or use the formula?

OpenStudy (amistre64):

becuase 10*10 is not equal to 10

OpenStudy (anonymous):

I don't understand how that relates...

OpenStudy (amistre64):

(x+2)(x-1) = 10 ... but 10 has many real factors so this setup is not all that useful. if we know when (x+2)(x-1) - 10 = 0 its more useful

OpenStudy (amistre64):

if (x+2) = 10, then x-1=1 ... and we would have x=2 2+2 is not 10

OpenStudy (amistre64):

if we have x-1 = 10, then x+2 = 1 and x=-2 but -2-1 is not 10

OpenStudy (amistre64):

lol, not x=-2 but same issue develops

OpenStudy (amistre64):

the issue is: 10*___ = 10 only if ___=1 the usefulness of equating stuff to zero is that: 0*____ = 0 is true for whatever ____ is

OpenStudy (anonymous):

aaaaaah!! ok, that makes more sense then :) So then, for normal parabola questions and such, I can simply break it up like so, but for logarithmic questions, I have to use the formula? Is that correct?

OpenStudy (amistre64):

for this particular case; the log can be simplified to a parabola question in order to find solutions

OpenStudy (anonymous):

Thank you very much! :D

OpenStudy (amistre64):

once you find the solutions, you will need to dbl chk that they make sense in the original setup. since log(a) only has a real value of a is greater than 0

OpenStudy (anonymous):

then since the answers to the formula were 3 and -4, 3 is the only possible answer. Got it :)